# “Would You Want an AI Tutor?” Understanding Stakeholder Perceptions of LLM-based Chatbots in the Classroom

### v1

@@ 3. A Review of perceptions of LLM-based chatbots
3.1. Methods
3.1.1. Selection procedure
We used Google Scholar to search for scientific literature on the topics of interest. To be considered, articles had to include the search strings (“Large language models” OR “LLMs” OR “LLM-based tutor” OR “LLM-powered chatbot” OR “ChatGPT”) AND (“perceptions” OR “perspectives” OR “attitudes”) in their title, abstract, keywords, or introduction. For the selection procedure, we limited the number of results to the first 100 articles, using Google Scholar’s relevance sort order. Articles were included if they met the following criteria: (1) discussed LLM-powered bots in education; (2) addressed perceptions of education stakeholders, such as students, teachers, parents, school staff, or education-related professionals; (3) were based on experimental, quasi-experimental, pre-experimental, or non-experimental research (excluding commentaries, secondary data analyses, and literature reviews); and (4) were written in English. Finally, the study must involve LLM-powered chatbots. Studies that only mentioned “AI tool” without specifying whether they referred to large language model-based technology or rule-based systems were excluded. This resulted in the selection of $N=38$ full-text articles for analysis.
3.1.2. Content analysis
3.2. Findings
3.2.1. Basic summary information
We now summarize the basic break-down of the $N=38$ articles included in our analysis. As shown in Figure 2(a), the majority of the included articles had samples from Asia ($N=21$), followed by Europe ($N=20$), North America ($N=18$), Oceania ($N=5$), Latin America ($N=4$), and Africa ($N=4$). (Note that several articles included samples from multiple continents.) When the location of the participants was not specified, the location of the authors’ university was selected as a proxy.
(a) continents
(b) educational contexts
Figure 2. Break-down of statistics of the analyzed articles ($N=38$).
Regarding the contexts of application of the LLM-based chatbots, Figure 2(b) illustrates that the most frequent educational context of the studies was STEM ($N=19$, 50%), followed by English-as-foreign-language (EFL) ($N=5$, 13%) and Humanities ($N=1$). A considerable number of studies did not specify the educational context ($N=13$, 34%).
3.2.2. RQ1: Key stakeholders and their perceptions
Figure 3. Stakeholders whose perceptions were analyzed in the articles ($N=38$).
@@ 3.2.2. RQ1: Key stakeholders and their perceptions
As shown in Figure 3, students were the most studied stakeholder group, appearing in 31 of the 38 (82%) studies, 27 of which involved University-level students and 4 involved K-12 students. Teacher perceptions were analyzed in 12 (32%) studies. Additionally, perceptions of industry professionals were considered in 1 study, albeit without explicitly specifying whether these were ed-tech professionals or learners. No studies included any other stakeholders, such as parents or school administrators.
@@
We identified different types of perception—positive (opportunities/benefits), negative (concerns), and neutral— in the articles we analyzed. We organized these perceptions into five categories: Attributes, Impact on Processes or Outcomes, Functions, User Experience, and Ethical and Societal Implications. Figure 4 summarizes results of our analysis, showing the total number of mentions of perceptions in each category separately for two stakeholder groups, Students (Figure 4(a)) and Teachers (Figure 4(b)). Complete results, with frequencies for each perception type within each category, are provided in Appendix A. We present additional details about each perception category in the remainder of this section. Note that, while we do not provide quantitative details about industry professionals as a stakeholder group in either Figure 4 or Appendix A, because they were only covered in one article, we do include them when discussing qualitative findings.
@@
Overall, the perceptions of both students and teachers were more frequently positive than either negative or neutral: with 58% of all perceptions were positive and 33% were negative for students vs. 51% positive and 38% negative for teachers. Additionally, while both students’ and teachers’ perceptions focused heavily on the ethics of LLM-based chatbots, particularly in a negative or concerned light, students also expressed views on UX, whereas teachers concentrated on the functions of LLM-powered bots. In contrast to the ethics of LLM-based bots, the topics of UX and functions were primarily addressed by their respective stakeholders in terms of opportunities or positive aspects.
@@
This category refers to perceptions related to teaching and learning processes and their results. These perceptions were most commonly expressed by students and teachers, and include: Efficiency, Student motivation, Performance expectancy (a concept from UTAUT (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), see Appendix B for a summary of these theories), referring to improved student learning and outcomes, Cost-effectiveness, as well as Critical thinking and problem solving skills. Among these, Performance expectancy, Cost-effectiveness, and Critical thinking and problem solving skills generated mixed perceptions. No neutral perceptions were found regarding these impacts.
@@
Figure 4. Summary of perceptions of two major stakeholders, students (4(a)) and teachers (4(b)), reported in the articles in our literature review. Here, Impact refers to “Impacts on Processes and Outcomes,” UX refers to “User Experience,” and Ethics refers to “Ethical and Societal Implications.” See Appendix A for a detailed break-down.
Functions.
This category refers to perceptions related to uses of LLM-powered technology in the classroom, including Assessment, Personalized tutoring, Material creation, Writing assistance, Brainstorming assistance, Opportunity for practice, Tool for complementing traditional teaching, Facilitating conditions (a construct in UTAUT that describes the degree to which an individual believes an organization’s and technical infrastructure support the use of the system (Venkatesh et al., 2003)), as well as Other functions. Some of these perceptions were viewed both positively, negatively, and neutrally, except for Tool for complementing traditional teaching, which was only perceived neutrally.
The most recurring positive perception about functions across all stakeholders was Personalized tutoring. Students recognized “the potential for personalized learning support” (Chan and Hu, 2023) and reported that “ChatGPT excels in breaking down intricate concepts into more understandable terms” (Zimotti et al., 2024). Teachers also emphasized the value of “Personalized teaching/24 hours Teaching Assistant (TA) access” (Ghimire et al., 2024), while industry professionals underscored that LLMs “support human effort” (Chen et al., 2024).
The most common negative perception regarding functions among students and teachers related to Facilitating conditions. For students, Ogbo-Gebhardt and Ogbo (2024) highlighted a lack of institutional support on GenAI usage (Chan and Hu, 2023; Ogbo-Gebhardt and Ogbo, 2024). For teachers, concerns included banning LLM-based tools (Lau and Guo, 2023) and the “lack of teacher support for using ChatGPT as an educational tool” (Iqbal et al., 2022). Similarly, the most commonly cited neutral perspective among students involved Facilitating conditions. For example, students in Woithe and Filipec (2023) stated, “[…] addressing these issues, not by banning ChatGPT, but by implementing regulatory measures and ethical guidelines for its use.” Teachers, on the other hand, emphasized the need for support in implementing LLM-based bots in the classroom. As one teacher in Iqbal et al. (2022) noted, “I feel like there should be more training available on how to use ChatGPT in the classroom. It would really help if we could get more help from the administration in this.”
@@
This category of perceptions relates to how individuals feel about interacting with LLM-powered bots. It includes concepts such as perceived Usefulness, Ease of use, Intention to use, Hedonic motivation, Well-being, Perceived importance, and Human interaction loss. Many of these—like Usefulness, Ease of use, Intention to use, and Hedonic motivation—originated in the Technology Acceptance Model (TAM) (Davis, 1989) and evolved in the Unified Theory of Acceptance and Use of Technology (UTAUT) (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), often with different names (see Appendices B and C). For instance, perceived Usefulness in TAM is called Performance expectancy in UTAUT (Sair and Danish, 2018), but we retained the term perceived Usefulness due to its use in studies beyond these frameworks. Similarly, TAM’s Ease of use evolved into UTAUT’s Effort expectancy, reflecting the role of technical skills in technology acceptance (Venkatesh et al., 2003). Except for Human interaction loss, which is perceived only negatively, and Hedonic motivation, viewed positively, negatively, and neutrally, all other concepts are perceived both positively and negatively.
@@
This category covers the anticipated impact of introducing LLM-based bots in schools, including Social justice, Job replacement, Reliance, Ethical and privacy concerns, Academic integrity, Social influence (a construct from UTAUT that captures the impact of family and friends on technology acceptance), Workplace changes, and Learning changes. While all were viewed positively and negatively, the last four also had neutral perceptions.
@@
3.2.3. RQ2: Gaps in the literature
@@
The second gap highlighted by our literature review was the inadequate representation of the educational context, particularly in terms of discipline and educational levels of the students’ sample. Figure 2(b) illustrates the contexts of application of LLM-based chatbots, with the most frequent ones being STEM, followed by studies with unspecified contexts. Additionally, a lack of diversity in educational levels could be observed in Figure 3, where 27 out of the 31 studies involving students focusing on university-level students and only four studies involving K-12 students. This is a critical issue because, while university students can independently choose whether to use AI technologies, K-12 students typically do not have a say in the implementation of these technology within their curricula (Singer, 2023; of Palm Beach County, 2024). This underscores the need for further research on perceptions across a wider range of educational contexts.
@@
Finally, the wide variety of theories, scales, and terms used across the literature to assess and name perceptions highlights the need for standardization in these aspects (see Appendices B and C). For example, in Ogbo-Gebhardt and Ogbo (2024), the theme of trust in and accuracy of the responses generated by LLM-based bots is considered a Facilitating condition (UTAUT), whereas in multiple other papers (Kazemitabaar et al., 2024; Ding et al., 2023; Yilmaz and Yilmaz, 2023) it is treated as a separate concept. This lack of unified terminology makes it difficult to develop a clear understanding of how stakeholders perceive LLM-based technology. We advocate for the standardization of terms in the field.
4. Proposed Framework
We propose a methodological framework, Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) (Figure 1), to guide policymakers, researchers, and developers in adopting AI-based chatbots more responsibly in educational contexts by incorporating stakeholder perceptions. Co-PACE addresses systematic gaps in the literature while supporting the iterative refinement of its nested taxonomy of perceptions.
The creation of Co-PACE is articulated around the perceptions of the different actors involved in the use of AI chatbots in education. Specifically, it is based on a stakeholder-first approach that aims to support key agents in designing and implementing transparent, compliant, and accountable AI systems (Bell et al., 2023). The second component is the goals that stakeholders have in mind when designing and/or implementing AI chatbots in educational settings. The context component was also added to address the lack of specification of the educational settings in which LLM-based chatbots are used, as shown in the literature review. And finally, the taxonomy of perceptions was added as a ’checklist’ for policymakers, technologists, and researchers to consult when assessing the presence of specific perceptions in a given setting and is designed to be refined over iterations through future studies. In the remainder of this section, we will describe each component in depth: stakeholders, context, goals, and taxonomy of perceptions, and explain how the framework should be used.
4.1. Stakeholders
We highlighted how the introduction of LLM-based chatbots, such as Khanmigo, in the classroom could be seen as a top-down decision-making process, where partnerships between private companies and governmental institutions or school districts often make decisions on behalf of direct stakeholders, such as teachers and students, or indirect stakeholders, such as parents and school staff. This echoes a concern already highlighted in the literature regarding how private companies, more often than community members and public servants, set pedagogy and policy in practice (Zeide, 2019). In contrast to this tendency, our framework prioritizes educational stakeholders: we identified students and teachers as the most recurring stakeholders in the literature; however, the systematic gaps we found underscore the need to assess the perceptions of other key stakeholders, such as parents, school staff, ed-tech professionals, and government agencies.
4.2. Context
As pointed out in Section 3.2.3, most studies in the literature have been conducted in the field of higher education. This underscores the need to assess the perceptions of stakeholders at other educational levels, such as younger students who lack the autonomy to make independent decisions, yet are still expected to use LLM-based chatbots as part of their curricula. The gaps also indicate that most studies focus on the STEM field, with many others having unspecified areas of application. We advocate for the need to assess perceptions of LLM-based chatbots across various disciplines.
4.3. Goals
The goals of soliciting and incorporating stakeholders’ perceptions when developing and deploying AI systems are diverse. In our framework, we propose goals drawn from the literature on responsible design of AI systems and organize them into two focus areas: those focused on system behavior and those focused on educational practice. When the focus is on system behavior, the goals are related to designing these tools or assessing whether they operate according to responsible AI key principles. We align with the existing literature on transparency, which asserts that the goal of a transparent design must begin with identifying the stakeholders (Bell et al., 2023; Chaudhry et al., 2022). Bell et al. (2023) established six categories to assist technologists in designing transparent, regulatory-compliant AI systems: validity, trust, learning and support, recourse, fairness, and privacy. Chaudhry et al. (2022) additionally emphasizes safety as a crucial component of transparency, especially in education, where mishaps in AI-powered technologies can often go unnoticed unless reported to the relevant authorities by a teacher or a student. Finally, although Louie et al. (2022) discuss this in relation to social robots rather than LLM-based chatbots, they distance themselves from the traditional designer-determined approach, advocating for the inclusion of culturally and linguistically diverse stakeholders as co-designers and expert informants in the design process. They promote participatory design, reporting that such inclusion leads to more engaging designs, improved user experience, and greater academic gains and ownership of technology-based learning.
In terms of pedagogical goals, the focus is on embedding the responsible principles in educational practices and relating them to specific learning outcomes (Domínguez Figaredo, 2020; Fu and Weng, 2024). For example, the in-depth analysis of Jurenka et al. (2024) refers to the specific learning practices embedded in the LearnLM tutor that the authors evaluated to improve the pedagogical capabilities of the AI system while promoting more responsible and effective learning experiences, including: evaluative practices, feedback on procedural problems, grounding of learning materials, adaptivity to learner level, engagement, etc. In Co-PACE, we have presented some of the goals that could be associated with implementing responsible practices with AI systems in education, and this is an area that practitioners could expand according to their vision and needs.
4.4. Taxonomy of perceptions
Our survey identified a gap in the inconsistent terminology used for similar concepts. To address this, our proposed taxonomy serves as a checklist for researchers, ed-tech professionals, and policymakers to assess, refine, and expand stakeholder perceptions, promoting greater consistency in perception research.
4.5. Putting the Co-PACE framework into practical use
The Co-PACE framework is designed to assist policy makers, researchers and technologists in two areas:
  1. (1)
    Elicit stakeholders’ perceptions to create new or refine existing recommendations to ensure that AI systems are human-centered and responsibly adopted.
  2. (2)
    Guide researchers and practitioners to address gaps in the literature on LLM-based chatbot perceptions through future studies.
In practice, the application of Co-PACE follows the sequence shown in Figure 1: starting with the first component, “stakeholders,” one should consider one or more of the bubbles before moving on to the next component. For example, policymakers and technologists of an ed-tech company might want to assess the perceptions of the parents of elementary school students of LLM-based chatbots when they are applied to humanistic subjects with the goal of improving or assessing their perceived effectiveness. Upon reaching the final component, the taxonomy, one could use it as a checklist to identify the presence of specific perceptions. In this example, when discussing effectiveness, perceptions related to Impact on Processes or Outcomes are likely to arise. Another example involves assessing teachers’ perceptions of LLM-based chatbots applied to English-as-a-Foreign-Language. The goals might include improving their experience in using this technology, as well as managing potential risks. When going through the taxonomy “checklist,” likely perceptions related to User Experience (UX) and Ethical and Societal Implications, such as cultural understanding (included in the “Social justice” perception), might arise. Co-PACE is an open framework, and so any missing elements in any component that emerge from conducting studies in less explored scenarios could be added.
5. Conclusion, Limitations, and Future Work
LLM-based chatbots are being piloted in schools, often through partnerships between private companies and school districts. However, feedback from direct users (students and teachers) and indirect users (parents and school staff) has been largely absent before and during rollout, aside from a few informal efforts. To address this gap, we conducted a literature review on stakeholders’ perceptions of LLM-based chatbots in education, identifying key positive, negative, and neutral views. Students generally perceive these chatbots as useful and efficient but raise concerns about accuracy. Teachers value their role as personalized tutors but are primarily concerned with maintaining academic integrity. Neutral perceptions from both groups describe chatbots as complementary to traditional teaching. The review also revealed gaps in the literature: (i) underrepresentation of key stakeholders, such as parents; (ii) limited exploration of diverse educational contexts, including various levels and disciplines; and (iii) inconsistent assessment of perceptions, with varying theories, scales, and terminology applied to similar concepts. We then proposed the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, incorporating a taxonomy of perceptions from the literature. Co-PACE is designed to help policymakers and technologists assess stakeholder perceptions throughout the development and deployment of AI systems in education, promoting more responsible and trustworthy processes.
Limitations
This article has limitations that must be considered when interpreting the findings. First, the organization of stakeholders’ perceptions into the taxonomy, as well as the assignment of values (positive, negative, or neutral) to them, was based on the researchers’ judgment during the content analysis process. Future iterations could help make the taxonomy more rigorous. Second, the literature review was limited to articles published on Google Scholar. Including articles from databases such as Scopus, ISI Web of Science, and other cross-references could add weight and rigor to the findings related to stakeholders’ perceptions of LLM-based chatbots in education. Lastly, the in-depth analysis of the literature was based on empirical studies conducted globally. Narrowing the scope to a particular region or tool could provide deeper insights, as perceptions may be influenced by cultural and social norms, socio-economic status, and other factors related to the context and background of the stakeholders.
Future work
There are several promising directions for future research. First, the stakeholder-first approach described here lays the foundation for future studies investigating neglected stakeholders’ perceptions across a range of overlooked contexts. Second, a repository of findings derived from the application of this framework could be created, to act as a reference for technologists and policymakers in future technology design processes and implementations.
Appendix A Supplementary statistics
@@
Complete break-down of perceptions, for all perception types in our taxonomy, for two main stakeholder groups: students and teachers. These results complement the summary in Figure 4 in the main body of the paper.
Appendix B Models of Perception

### v2

@@ A.2. Findings
A.2.1. RQ: Key stakeholders and their perceptions
@@
As shown in Figure 2 (right), students were the most studied stakeholder group, appearing in 31 of the 38 (82%) studies, 27 of which involved University-level students and 4 involved K-12 students. Teacher perceptions were analyzed in 12 (32%) studies. Additionally, perceptions of industry professionals were considered in 1 study, albeit without explicitly specifying whether these were ed-tech professionals or learners. No studies included any other stakeholders, such as parents or school administrators.
We identified different types of perceptions: positive (opportunities/benefits), negative (concerns), and neutral. We organized these perceptions into four categories: Attributes, Impact on Processes or Outcomes, User Experience, and Ethical and Societal Implications.
@@
Figure 3 summarizes results of our analysis, showing the total number of mentions of perceptions in each category separately for two stakeholder groups, Students (Figure 3(a)) and Teachers (Figure 3(b)). We present additional details about each perception category in the remainder of this section. Note that, while we do not provide quantitative details about industry professionals as a stakeholder group, because they were only covered in one article, we do include them when discussing qualitative findings.
@@
Overall, the perceptions of students were more frequently positive than either negative or neutral, whereas teachers’ were slightly more frequently negative. Specifically, with 56% of all perceptions were positive and 38% were negative for students vs. 43% positive and 52% negative for teachers. Additionally, while both students’ and teachers’ perceptions focused heavily on the ethics of LLM-based systems, particularly in a negative or concerned light, students also expressed multiple views on UX, particularly emphasizing opportunities and positive aspects. In contrast, teachers expressed fewer opinions on UX as well as on other dimensions.
@@
This category refers to perceptions related to teaching and learning processes and their results. These perceptions were most commonly expressed by students and teachers, and include: Efficiency, Student motivation, Performance expectancy (a concept from UTAUT (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), referring to improved student learning and outcomes, Cost-effectiveness, as well as Critical thinking and problem solving skills. Among these, Performance expectancy, Cost-effectiveness, and Critical thinking and problem solving skills generated mixed perceptions. No neutral perceptions were found regarding these impacts.
@@
Figure 3. [Summary of perceptions of two major stakeholders, students (3(a)) and teachers (3(b)), reported in the articles in our literature review. Here, Impact refers to “Impacts on Processes and Outcomes,” UX refers to “User Experience,” and Ethics refers to “Ethical and Societal Implications.” See Appendix B for a detailed break-down.
@@
This category of perceptions relates to how individuals feel about interacting with LLM-powered systems. It includes concepts such as perceived Usefulness, Ease of use, Intention to use, Hedonic motivation, Well-being, Perceived importance, and Human interaction loss. Many of these—like Usefulness, Ease of use, Intention to use, and Hedonic motivation—originated in the Technology Acceptance Model (TAM) (Davis, 1989) and evolved in the Unified Theory of Acceptance and Use of Technology (UTAUT) (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), often with different names. For instance, perceived Usefulness in TAM is called Performance expectancy in UTAUT (Sair and Danish, 2018), but we retained the term perceived Usefulness due to its use in studies beyond these frameworks. Similarly, TAM’s Ease of use evolved into UTAUT’s Effort expectancy, reflecting the role of technical skills in technology acceptance (Venkatesh et al., 2003). Except for Human interaction loss, which is perceived only negatively, and Hedonic motivation, viewed positively, negatively, and neutrally, all other concepts are perceived both positively and negatively.
@@
This category covers the anticipated impact of introducing LLM-based tools in schools, including Social justice, Job replacement, Reliance, Ethical and privacy concerns, Academic integrity, Social influence (a construct from UTAUT that captures the impact of family and friends on technology acceptance), Workplace changes, and Learning changes. While all were viewed positively and negatively, the last four also had neutral perceptions.
@@
A.3. Gaps in the literature
After reviewing relevant academic work on stakeholder perceptions of AI-based systems in educational settings, we identified the following gaps:
@@ A.3. Gaps in the literature
The second gap highlighted by our literature review was the inadequate representation of the educational context, particularly in terms of discipline and educational levels of the students’ sample. Figure 2 illustrates the contexts of application of LLM-based systems, with the most frequent ones being STEM, followed by studies with unspecified contexts. Additionally, a lack of diversity in educational levels could be observed in Figure 2, where 27 out of the 31 studies involving students focusing on university-level students and only four studies involving K-12 students. This is a critical issue because, while university students can independently choose whether to use AI technologies, K-12 students typically do not have a say in the implementation of these technology within their curricula (Singer, 2023; of Palm Beach County, 2024). This underscores the need for further research on perceptions across a wider range of educational contexts.
@@
Finally, the wide variety of theories, scales, and terms used across the literature to assess and name perceptions highlights the need for standardization in these aspects. For example, in Ogbo-Gebhardt and Ogbo (2024), the theme of trust in and accuracy of the responses generated by LLM-based bots is considered a Facilitating condition (UTAUT), whereas in multiple other papers (Kazemitabaar et al., 2024; Ding et al., 2023; Yilmaz and Yilmaz, 2023) it is treated as a separate concept. This lack of unified terminology makes it difficult to develop a clear understanding of how stakeholders perceive LLM-based technology.
Building on the works described above, we advocate for the inclusion and systematic collection of stakeholder perspectives to inform the design and deployment of LLM-powered systems in the classroom.
Appendix B Supplementary statistics
@@
Complete break-down of perceptions, for all perception types in our taxonomy, for two main stakeholder groups—students and teachers—are presented in Table 6.
Table 6. Frequency of positive, negative, and neutral perceptions of AI tutors by students and teachers
Perception Students (Positive) Teachers (Positive) Students (Negative) Teachers (Negative) Students (Neutral) Teachers (Neutral)
Attributes
Convenience 16 3 0 0 0 0
Accuracy 14 2 20 6 0 0
Knowledge base 7 2 4 0 0 0
Anthropomorphism 9 1 4 1 0 0
Other technical specs 3 0 2 2 0 0
Total (Attributes) 49 8 30 9 0 0
Impact on Processes / Outcomes
Efficiency 17 6 0 0 0 0
Performance expectancy 11 1 1 0 0 0
Student motivation 5 2 0 0 0 0
Critical thinking & problem solving 6 2 16 6 0 0
Cost-effectiveness 0 1 0 1 0 0
Total (Impact) 39 12 17 7 0 0
User Experience
Ease of use 16 4 3 2 0 0
Intention to use 15 2 3 1 0 0
Usefulness 19 3 4 0 0 0
Hedonic motivation 12 1 5 1 1 0
Well-being 8 1 2 0 0 0
Perceived importance 1 0 0 0 0 0
Human interaction loss 0 0 5 2 0 0
Total (UX) 71 11 22 6 1 0
Ethical & Societal Implications
Social justice 6 4 6 5 0 0
Job replacement 3 2 8 3 0 0
Reliance 8 0 14 6 0 0
Ethical/privacy concerns 2 1 12 4 0 0
Academic integrity 3 1 12 8 3 0
Social influence 4 1 5 0 1 0
Changes in workplace 0 0 0 0 9 2
Changes in learning 0 0 0 0 4 2
Total (Ethics) 26 9 57 26 17 4
Grand Total 185 40 126 48 18 4
Table 7. Co-PALE mapping for Case 2: AI Curriculum Support Tool in High School
Stakeholder Context R-AI Goals Educational Goals
Teachers Pilot implementation in high school curriculum (grades 9–12) Fairness in grading, transparency, data security Improve instructional efficiency; support differentiated learning
Perceptions
Attributes. Accuracy and reliability of grading are essential. User Experience. Ease of use impacts integration into curriculum. Impact on Process / Outcome. Time savings and teaching support noted. Ethical and Societal Implications. Fairness and data privacy are key concerns.
Students Direct users of the tool’s outputs (assignments, grades, feedback) Privacy, fairness, and autonomy in AI-supported evaluation Receive effective, personalized, and fair instruction
Perceptions
Attributes. Accuracy and fairness in grading matter deeply. User Experience. Clarity and usefulness of feedback are important. Impact on Process / Outcome. Influences motivation and engagement. Ethical and Societal Implications. Privacy and transparency are top concerns.
Parents Indirect stakeholders in secondary education Transparency in data use; fairness in learning outcomes Ensure academic quality; protect student well-being
Perceptions
Attributes. Expect transparency in system function. User Experience. Usability contributes to trust. Impact on Process / Outcome. Concerned with grading fairness. Ethical and Societal Implications. Strong concern for data use and fairness.
School Administrators Oversight of AI integration into school-wide curriculum delivery Compliance, equity, transparency, and oversight responsibilities Support instructional innovation; manage resources effectively
Perceptions
Attributes. Focus on implementation and reliability. User Experience. Relevant to monitoring usage and uptake. Impact on Process / Outcome. Links to teacher performance and student outcomes. Ethical and Societal Implications. Oversight of fairness and equity in deployment.
Ed-tech Professionals Design and deployment of curriculum support systems Accuracy, fairness, and technical compliance Enable scalable solutions for personalized learning
Perceptions
Attributes. Accuracy and specification requirements are critical. User Experience. Interface usability informs deployment success. Impact on Process / Outcome. Efficiency influences implementation. Ethical and Societal Implications. Compliance prioritized over broader ethics.
Appendix C Models of Perception